OpenStudy (anonymous):

Determine whether each situation involves a permutation or a combination. Then find the number of possibilities 1). 8 cars in a row parked next to a curb 2). a hand of 6 cards from a standard deck of cards

OpenStudy (amistre64):

permutation sees different orders as different elements combinations see sets as an element so order doesnt matter

OpenStudy (amistre64):

8 cars can be parked in different ways; 8! ways 6 cards from a standard deck are the same regardless of how you hold them: P(52,6),6!

OpenStudy (anonymous):

2) 52!/(52-6)!*6! ?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

1). 8 cars in a row parked next to a curb here arrangement is considered so it's permutation 8P1=8 2). a hand of 6 cards from a standard deck of cards here it's a cobination as no arrangements is considered 52C6=

OpenStudy (amistre64):

C(n,r) is just P(n,r)/r!

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

\[\dbinom{52}{6}=\frac{52\times51\times 50\times 49\times 48\times 47}{6\times5 \times 4\times 3\times 2}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

this is of course a whole number, so cancel everything before you multiply

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\dbinom{52}{6}=\frac{52\times51\times 50\times 49\times 48\times 47}{6\times5 \times 4\times 3\times 2}=52\times 17\times 10\times 49\times 47\]